OpenStudy (anonymous):

@Mathematics @Mathematics @Mathematics 3 students are selected at random from a class consisting of 5 freshman, 3 sophomores, and 2 juniors. Find the probability that 2 freshman and 1 junior are selected given that at least 1 freshman is selected.

OpenStudy (zarkon):

2/11

OpenStudy (anonymous):

please show me. i can find probability of choosing 2 freshman and 1 junior, that is 5C2 * 2C1 / ( 10 C 3 )

OpenStudy (zarkon):

P(2f,j|at least 1 f)=P(2f,j,at least 1 f)/P(at least 1 f) =P(2f,j)/P(at least 1 f)=P(2f,j)/(1-P(0 f))

OpenStudy (anonymous):

wait, i dont follow

OpenStudy (anonymous):

ok i see what you did in the denominator,

OpenStudy (anonymous):

P(2f,j,at least 1 f) = P(2f,j) ?

OpenStudy (zarkon):

the event {2f} is a subset of {at least one f} so the intersection is {2f}

OpenStudy (anonymous):

ok, so P ( 2f j & at least 1 f ) = P ( 2f j )

OpenStudy (anonymous):

but its not a subset

OpenStudy (zarkon):

it is

OpenStudy (anonymous):

because you could have 3,4,.. freshman

OpenStudy (zarkon):

{at least 1f}= {1f or 2f or 3f} since you are picking 3

OpenStudy (anonymous):

oh it is , { f1 , f2} subset of { f1,f2,f3,f4,f5 } ?

OpenStudy (anonymous):

, where we dont identify f1,...

OpenStudy (zarkon):

{2f} intersect {1f or 2f or 3f} ={2f}

OpenStudy (anonymous):

there are 5 freshman actually

OpenStudy (zarkon):

you are only picking 3 people though

OpenStudy (zarkon):

so the most freshmen you can pick is 3

OpenStudy (anonymous):

but, you could pick one freshman exactly, in which case {2f} is not subset of {1f}

OpenStudy (zarkon):

no...you are not thinking of it in the right way

OpenStudy (zarkon):

the event is {1f or 2f or 3f}

OpenStudy (anonymous):

ok the whole event

OpenStudy (anonymous):

alright, can you do this in a tree diagram ?

OpenStudy (zarkon):

it is discrete so you can

OpenStudy (anonymous):

by the way, you left out operation of and or

OpenStudy (zarkon):

?

OpenStudy (anonymous):

P(1f &2f & j|at least 1 f)=P(2f& j&at least 1 f)/P(at least 1 f) etc =P(2f,j)/P(at least 1 f)=P(2f,j)/(1-P(0 f))

OpenStudy (anonymous):

you could abbreviate and say 2f = 1f & 2f

OpenStudy (zarkon):

no

OpenStudy (anonymous):

P(1f &2f &1 j|at least 1 f)=P(1f&2f&1 j&at least 1 f)/P(at least 1 f)

OpenStudy (zarkon):

2f and 1f & 2f are two different things

OpenStudy (anonymous):

ok what did you mean by 2f ?

OpenStudy (zarkon):

1f & 2f can't happen

OpenStudy (zarkon):

that you picked 2 freshmen

OpenStudy (anonymous):

1f = choice of first freshmen, or first draw you picked a freshman and 2f = second choice of freshman, or second draw you picked a freshman

OpenStudy (zarkon):

I wouldn't use that notation...I would use \[f_1, f_2\]

OpenStudy (anonymous):

you used it before {2f} intersect {1f or 2f or 3f} ={2f}

OpenStudy (zarkon):

since we don't care about order anyway...I wouldn't pick them one at a time

OpenStudy (zarkon):

no...{2f}= you picked 2 freshmen {1f or 2f or 3f}= you pick 1 freshmen or 2 freshmen or 3 freshmen

OpenStudy (anonymous):

so you wouldnt do , first freshman drawn, then second freshman drawn on second draw

OpenStudy (zarkon):

no

OpenStudy (anonymous):

im thinking of a coin? where do we use drawing, oh cards

OpenStudy (anonymous):

first draw a heart , then draw a spade. that sort of thing

OpenStudy (anonymous):

but then you have the issue of order does not count , so it makes it complicated

OpenStudy (zarkon):

that is if you care about the order...here you don't

OpenStudy (anonymous):

you have f f j, f j f , j f f

OpenStudy (anonymous):

if you break this problem into drawings

OpenStudy (zarkon):

sure

OpenStudy (anonymous):

but you didnt do it that way

OpenStudy (anonymous):

or, i mean selections

OpenStudy (zarkon):

you can...I didn't

OpenStudy (anonymous):

one selection at a time

OpenStudy (anonymous):

ok so you did P ( 2f & 1j | at least 1j ) / ( P ( at least 1 j )

OpenStudy (zarkon):

close

OpenStudy (anonymous):

you wrote P ( 2f, j) , the comma threw me off

OpenStudy (anonymous):

what does comma mean exactly ?

OpenStudy (zarkon):

and

OpenStudy (zarkon):

2f,j=2f & j

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but 2f = { 1f or 2f } , no it doesnt

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

and P ( 2f, j) = 5C2 * 2C1 / ( 10 C3 ) ?

OpenStudy (zarkon):

\[2f\neq1f\text{ or }2f\] using my notation

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

and how did you get P ( at least 1f) = 1 - P ( 0 F)

OpenStudy (zarkon):

complement

OpenStudy (anonymous):

since the event is choosing 3 , then we have , i dont know

OpenStudy (anonymous):

what is P( 0f)

OpenStudy (zarkon):

0f=pick no freshmen

OpenStudy (anonymous):

the three students selected can be sophomores or juniors >

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

you can lump them together

OpenStudy (anonymous):

so it will be 5C3 / 10 c 3

OpenStudy (anonymous):

errr, i dont know , the numerator is wrong

OpenStudy (zarkon):

\[\frac{\frac{_5C_2\cdot_2C_1}{_{10}C_3}}{1-\frac{_5C_3}{_{10}C_3}}\]

OpenStudy (anonymous):

so i had P(0f) right

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

but if there are no freshman, dont we have to consider the sophomores and juniors seperately

OpenStudy (zarkon):

we can...but it is quicker to temporally label them collectively as non-freshmen

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and 10 C 3 because thats the whole event

OpenStudy (zarkon):

yes...that is the experiment ...picking 3 from 10 where order does not matter

OpenStudy (anonymous):

how do you know that 2f,j = 2f & j, sometimes ive seen it as or

OpenStudy (zarkon):

a , should be treated as & when inside a probability

OpenStudy (anonymous):

not as an or. that would make the problem different if it was an 'or'

OpenStudy (zarkon):

you can always use and if you want

OpenStudy (zarkon):

yes...that would change the problem if it was an 'or'

OpenStudy (anonymous):

hmmm, what would it be? and can we do this problem over with a tree model or tree

OpenStudy (anonymous):